# CaTS-Bench: Can Language Models Describe Numeric Time Series?

### v1

@@ H.4.3 Latent Similarities
Table 13: Embedding similarities
Encoder Model Dimensionality Mean Median STD Intra Inter % Pairs >0.95>0.95
MiniLM-L12-v2 368 0.2932 0.2500 0.1771 0.6202 0.2269 5.26
MiniLM-L6-v2 384 0.3339 0.3013 0.1510 0.5999 0.2800 1.47
mpnet-base-v2 768 0.3278 0.2791 0.1775 0.6631 0.2599 8.29
bge-large-v1.5 1024 0.5807 0.5605 0.0986 0.7600 0.5443 1.67
mxbai-large-v1 1024 0.5293 0.5032 0.1116 0.7370 0.4871 2.15
Qwen3-4B 2560 0.3816 0.3530 0.1273 0.6098 0.3354 0.20
e5-mistral-7b 4096 0.6507 0.6358 0.0697 0.7770 0.6251 1.33
GritLM-7B 4096 0.4767 0.4537 0.0947 0.6481 0.4420 0.04
Qwen3-8B 4096 0.3736 0.3465 0.1225 0.5947 0.3288 0.32
Average - 0.4386 0.4092 0.1256 0.6678 0.3922 2.30
H.4.4 Human vs. AI Embedding Similarity
We computed cosine similarities between captions authored by humans, GPT-4o, and Gemini using Qwen-8B embeddings. Both small-scale ($22$ vs. $22$ captions) and larger-scale comparisons ($22$ vs. $150$ captions; $3k$ vs. $3k$ STOCK dataset (Jhamtani & Berg-Kirkpatrick, 2021)) show only modest gaps between human–AI and human–human similarity. For random CaTS-Bench samples, we report averages across five runs with deviations in the range 0.00100.00100.01740.0174. See Table 14 for full results.
Table 14: Cosine similarities between human, GPT-4o, and Gemini captions. Human–AI similarity closely tracks human–human similarity, suggesting minimal stylistic divergence.
Balanced (22 vs 22) Unbalanced (22 vs 150) STOCK (3k vs 3k)
Comparison Human GPT Gemini Human GPT Gemini Human GPT Gemini
Human 0.371 0.341 0.353 0.371 0.358 0.353 0.654 0.662 0.586
GPT 0.406 0.328 0.401 0.329 0.694 0.604
Gemini 0.370 0.378 0.599
H.5 Reproducibility Validation
We re-ran our evaluation three times on $\sim$500 randomly sampled test examples from the semi-synthetic set across five representative models (GPT-4o, Claude 3 Haiku, LLaMA, Idefics, Qwen-VL). Figure 6 shows a log-scale visualization. Across nearly all metrics, the variance is vanishingly small, often on the order of $10^{-6}$, which confirms stability and supports single-run robustness.
@@ H.5 Reproducibility Validation
Appendix I Role of Vision
I.1 Visual Modality Ablation
We present Table 15, comparing the performance of VLMs with and without the visual input.
Table 15: Evaluation of generated semi-synthetic captions under modality ablation. Each metric is split into two columns: VL (vision-language input) and L (text-only input).
Model DeBERTa F1 SimCSE BLEU ROUGE-L METEOR Numeric
VL L VL L VL L VL L VL L VL L
InternVL 0.659 0.677 0.794 0.854 0.081 0.113 0.247 0.283 0.260 0.317 0.589 0.636
QwenVL 0.643 0.648 0.790 0.802 0.081 0.090 0.249 0.258 0.260 0.274 0.503 0.520
Phi-4 M.I. 0.624 0.674 0.797 0.845 0.074 0.118 0.274 0.290 0.239 0.306 0.583 0.634
SmolVLM 0.594 0.603 0.692 0.758 0.043 0.068 0.224 0.244 0.178 0.257 0.473 0.565
Llama 3.2 Vision 0.670 0.669 0.850 0.849 0.117 0.110 0.290 0.275 0.314 0.313 0.684 0.598
Idefics2-8B 0.604 0.632 0.698 0.816 0.040 0.080 0.225 0.258 0.161 0.270 0.454 0.585
LLaVA v1.6 Mistral 0.650 0.648 0.820 0.824 0.086 0.098 0.259 0.261 0.287 0.286 0.517 0.532
Claude 3 Haiku 0.682 0.676 0.856 0.853 0.112 0.117 0.291 0.287 0.300 0.298 0.628 0.609
Gemini 2.0 Flash 0.688 0.698 0.858 0.871 0.137 0.175 0.318 0.343 0.279 0.328 0.677 0.684
I.2 Visual Attention Analysis
Interpreting visual grounding in large multimodal models is non-trivial, as not all of them expose interpretable cross-modal attention mechanisms. We attempt this using the LLaVA model, which provides access to decoder-level cross-attention weights over vision tokens. We adapt the approach in Zhang (2024) for the LLaVA 1.6 model. We visualize per-token visual grounding via the following steps. For each generated token, we extract the decoder cross-attention matrix $\mathbf{A}_{\text{llm}}\in\mathbb{R}^{T\times V}$, where $T$ is the number of generated tokens and $V$ is the number of vision tokens.
@@ I.2 Visual Attention Analysis
I.3 Testing Alternative Visual Encodings
@@
Table 16: Effect of alternative visual encodings on captioning performance on subset of semi-synthetic captions. Values show the baseline score for No Plot (TS+Text) and relative differences ($\Delta$) for line plots, Gramian Angular Fields (GAFs), and recurrence plots (RPs).
@@
   Idefics2 (8B)    DeBERTa F1    0.6255    -0.078    -0.0525    -0.0525
   BLEU    0.069    -0.0405    -0.0460    -0.0430
@@
Table 17: Model accuracy for time-series Q&A tasks. Bolded and underlined scores respectively denote first and second places (excluding human performance). Caption/Plot/TS refer to caption, plot, and time series matching. Amplitude/Peak Earlier/Mean/Variance refer to amplitude, peak, mean, and variance comparison. Green and Red indicate improvement and degradation after finetuning, respectively.
Model Caption Plot TS Amplitude Peak Earlier Mean Variance
@@
Gemini 2.0 Flash 0.78 0.30 0.61 0.8 0.42 0.7 0.62
Gemini 2.5 Pro Preview 0.66 0.30 0.31 1.0 1.0 1.0 0.85
Claude 3 Haiku 0.68 0.29 0.57 0.65 0.40 0.53 0.33
GPT-4o 0.96 0.31 0.77 0.825 0.725 0.7 0.625
@@
LLaVA v1.6 Mistral 0.39 0.27 0.32 0.45 0.45 0.42 0.45
Phi-4 M.I. 0.62 0.29 0.45 0.7 0.82 0.68 0.7
Idefics 2 0.49 0.25 0.29 0.35 0.4 0.4 0.5
SmolVLM 0.26 0.34 0.28 0.4 0.48 0.44 0.6
QwenVL 0.68 0.27 0.61 0.7 0.5 0.6 0.4
Llama 3.2 Vision 0.66 0.24 0.27 0.45 0.63 0.43 0.3
Finetuned
LLaVA v1.6 Mistral 0.44 0.25 0.29 0.43 0.53 0.35 0.4
Phi-4 M.I. 0.59 0.29 0.45 0.83 0.88 0.7 0.55
Idefics 2 0.33 0.23 0.29 0.58 0.38 0.5 0.63
SmolVLM 0.18 0.26 0.29 0.28 0.48 0.38 0.58
QwenVL 0.55 0.25 0.43 0.7 0.4 0.58 0.58
Llama 3.2 Vision 0.66 0.24 0.27 0.4 0.6 0.43 0.33
Human 0.81 0.95 0.83 0.925 0.85 0.95 0.90
J.1 Sample Q&A Questions
@@
We provide examples of Q&A questions in Figures 91012111314, and 15, covering one example per question type.
@@
Question
Here is a time series caption:
Answer
"answer": "D"
Figure 15: Example of a time series matching question.
J.2 Qwen-Based Filtering
@@
J.3 Distractor Generation in Q&A Tasks
@@
Time Series Matching Cross-domain 0.9803
Time Series Matching Same-domain 0.9586
Time Series Matching Artificially Perturbed 0.6864
@@
Appendix K Inference Examples
K.1 Numeric Time Series Ablation
@@
K.2 Statistical Inference Failure & Success
@@
Appendix L CaTS-Bench Samples
In this section, we provide representative time series samples from CaTS-Bench across several domains. Samples include the numeric time series segment, rich JSON metadata, line plot image, and ground-truth caption.
Time Series Segment
Metadata JSON
Line Plot Image
Caption
Time Series Segment
[29, 33, 29, 35, 32, 35, 36, 34, 23, 34, 26, 25, 24, 30, 31, 39, 34, 48, 56, 40, 35, 37, 42, 55, 57, 41, 47, 53, 56, 50, 59, 45, 49, 44, 36]
Metadata JSON
{ "border": "US-Canada Border", "end date of the series": "2022-11-01", "general maximum in the history of this port": 80, "general mean in the history of this port": 30.85, "general minimum in the history of this port": 0, "general standard deviation in the history of this port": 13.74, "maximum in this specific series": 59, "mean of this specific series": 39.4, "means": "Trains", "minimum in this specific series": 23, "port": "Van Buren", "sampling frequency": "monthly", "standard deviation of this specific series": 10.16, "start date of the series": "2019-12-01", "state": "Maine" }
Line Plot Image
Caption
From December 2019 to December 2022, the number of monthly trains crossing the port of Van Buren generally increased, with an average of 39.4 trains, which is notably higher than the all-time mean of 30.85. The time series starts with values in the range of 23 to 36 trains until November 2020, and then experiences a significant upward shift, reaching a peak of 59 trains in October 2022. This increase suggests a moderate surge in cross-border traffic compared to historical trends.
Time Series Segment
[0.52, 0.32, 0.30, 0.38, 0.41, 0.51, 0.43, 0.41, 0.47]
Metadata JSON
{ "attribute": "co2_emissions",
"country": "Djibouti",
"end year of this series": 2018,
"maximum of this specific series": 0.52,
"mean of this specific series": 0.42,
"minimum of this specific series": 0.30,
"population at the end year": 1071886.0,
"population at the start year": 930251.0,
"region": "Middle East & North Africa",
"sampling frequency": "yearly",
"standard deviation of this specific series": 0.07,
"start year of this series": 2010 }
Line Plot Image
Caption
Time Series Segment
[ 90.0, 90.0, 104.0, 104.0, 104.0, 104.0, 104.0, 104.0, 104.0, 110.57, 110.57, 110.57, 110.57, 110.57, 110.57, 110.57, 121.29, 121.29, 121.29, ..., 268.86, 268.86, 257.29, 257.29, 257.29, 257.29, 257.29, 257.29, 257.29, 266.14, 266.14, 266.14, 266.14, 266.14, 266.14, 266.14, 394.57, 394.57, 394.57, 394.57, 394.57, 394.57, 394.57, 411.57, 411.57, 411.57, 411.57, ..., 196.71, 196.71, 196.71, 143.43, 143.43, 143.43, 143.43, 143.43, 143.43, 143.43, 152.43 ]
Metadata JSON
Line Plot Image
Caption
Time Series Segment
[872817.62, 972716.24, 936508.43, 806979.15, 982322.24, 861894.77, 861941.25, 872469.03, 807798.73, 833517.19, 868191.05, 777207.3, 865924.2]
Metadata JSON
{ "attribute": "weekly_sales", "best week": "2010-12-24", "best week sales": 1309226.79, "end week of this series": "2012-08-31", "maximum of this specific series": 982322.24, "mean of this specific series": 870791.32, "mean sales": 893581.39, "minimum of this specific series": 777207.3, "sampling frequency": "weekly", "start week of this series": "2010-05-28", "worst week": "2010-12-31", "worst week sales": 635862.55 }
Line Plot Image
Caption
The provided sales data, spanning from May 2010 to August 2012, reveals fluctuations around a mean of \$870,791.32, which is moderately below the overall store mean of \$893,581.39. The lowest sales week within this period reached \$777,207.3, while the highest peaked at \$982,322.24, indicating a moderate level of volatility, but no extreme values compared to the store’s best and worst weeks. The absence of a clear upward or downward trend suggests that sales remained relatively stable during this specific timeframe.
Appendix M Editing and Review
M.1 Edits made in the captions
@@
M.2 Interface
@@
Appendix N Template-based Prompts
@@
N.1 Ground-Truth Caption Generation Prompt
@@
N.2 Baseline Caption Generation Prompt
@@
<time series>
Describe this time series by focusing on trends and patterns. Discuss concrete numbers you see and pay attention to the dates. For numeric values, ensure consistency with the provided time series. If making percentage comparisons, round to the nearest whole number. Report the dates when things happened.
Use descriptive language to create engaging, natural-sounding text. Avoid repetitive phrasing and overused expressions.
N.3 Caption Paraphrasing Prompt
@@
The full TSC prompt from N.2 is injected as the caption_prompt string.
N.5 Semantic Perturbation Prompt
@@
N.6 Numeric Perturbation Prompt

### v2

@@ G.4.3 Latent Similarities
Table 16: Embedding similarity statistics for human-written and synthetic captions across encoder models.
Encoder Dim Human Captions Synthetic Captions
Mean Med STD Intra Inter %>0.95>0.95 Mean Med STD Intra Inter %>0.95>0.95
MiniLM-L12-v2 384 0.2755 0.2437 0.1560 0.6090 0.2308 0.17 0.2932 0.2500 0.1771 0.6202 0.2269 5.26
MiniLM-L6-v2 384 0.3367 0.3113 0.1429 0.6291 0.2975 0.12 0.3339 0.3013 0.1510 0.5999 0.2800 1.47
mpnet-base-v2 768 0.3260 0.2859 0.1679 0.6866 0.2776 0.44 0.3278 0.2791 0.1775 0.6631 0.2599 8.29
bge-large-v1.5 1024 0.5921 0.5753 0.0955 0.7903 0.5655 0.23 0.5807 0.5605 0.0986 0.7600 0.5443 1.67
mxbai-large-v1 1024 0.5344 0.5111 0.1096 0.7691 0.5030 0.24 0.5293 0.5032 0.1116 0.7370 0.4871 2.15
Qwen3-4B 2560 0.3833 0.3613 0.1195 0.6344 0.3496 0.02 0.3816 0.3530 0.1273 0.6098 0.3354 0.20
Qwen3-8B 4096 0.3800 0.3571 0.1188 0.6306 0.3465 0.00 0.3736 0.3465 0.1225 0.5947 0.3288 0.32
e5-mistral-7b 4096 0.6469 0.6354 0.0670 0.7839 0.6285 0.01 0.6507 0.6358 0.0697 0.7770 0.6251 1.33
Average 0.4344 0.4101 0.1222 0.6910 0.3999 0.15 0.4386 0.4092 0.1256 0.6678 0.3922 2.30
G.4.4 Human vs. AI Embedding Similarity
We measure cosine similarity between captions authored by humans, GPT-4o, and Gemini using Qwen-8B embeddings to assess whether AI-generated captions diverge stylistically from human-written references. Across both CaTS-Bench and the STOCK dataset, we observe only modest gaps between human-AI and human-human similarity in embedding space, indicating overlap in semantic space despite substantial surface-level differences. This suggests that while AI captions differ lexically and structurally from human captions, they do not collapse to a distinct semantic mode.
For CaTS-Bench, we report results on a balanced subset of $500$ human and $500$ AI captions, averaged over five random draws. For STOCK, we evaluate $3k$ human and $3k$ AI captions following prior work. Full results are shown in Table 17.
Table 17: Cosine similarities between human, GPT-4o, and Gemini captions using Qwen-8B embeddings. Human-AI similarity is consistently lower than human-human similarity, but remains semantically comparable.
Comparison Human GPT Gemini
CaTS-Bench (500 vs 500)
Human 0.389 0.341 0.336
GPT 0.414 0.321
Gemini 0.372
STOCK (3k vs 3k)
Human 0.658 0.621 0.603
GPT 0.692 0.588
Gemini 0.601
G.5 Reproducibility Validation
We re-ran our evaluation three times on $\sim$500 randomly sampled test examples from the synthetic data across five representative models as generators (GPT-4o, Claude 3 Haiku, LLaMA, Idefics, Qwen-VL). Figure 8 shows a log-scale visualization. Across nearly all metrics, the variance is vanishingly small, often on the order of $10^{-6}$, which confirms stability and supports single-run robustness.
@@ G.5 Reproducibility Validation
Appendix H Role of Vision
H.1 Visual Attention Analysis
Interpreting visual grounding in large multimodal models is non-trivial, as not all of them expose interpretable cross-modal attention mechanisms. We attempt this using the LLaVA model, which provides access to decoder-level cross-attention weights over vision tokens. We adapt the approach in vlm_visualizer for the LLaVA 1.6 model. We visualize per-token visual grounding via the following steps. For each generated token, we extract the decoder cross-attention matrix $\mathbf{A}_{\text{llm}}\in\mathbb{R}^{T\times V}$, where $T$ is the number of generated tokens and $V$ is the number of vision tokens.
@@ H.1 Visual Attention Analysis
H.2 Testing Alternative Visual Encodings
@@
Table 18: Effect of alternative visual encodings on captioning performance on a subset of synthetic captions. Values show the baseline score for No Plot (TS+Text) and relative differences ($\Delta$) for line plots, Gramian Angular Fields (GAFs), and recurrence plots (RPs).
@@
Idefics2 (8B) DeBERTa F1 0.6255 -0.0780 -0.0525 -0.0525
BLEU 0.0690 -0.0405 -0.0460 -0.0430
@@ Appendix I Q&A Tasks
Table 19: Model accuracy for time-series Q&A tasks. Bold and underlined scores denote first and second places (excluding human). Cap/Plt/TS refer to caption, plot, and time-series matching. Amp/Peak/Mean/Var denote amplitude, peak-earlier, mean, and variance comparison.
Model Cap Plt TS Amp Peak Mean Var
@@
Gemini2 F 0.78 0.30 0.61 0.80 0.42 0.70 0.62
Gemini2.5 0.66 0.30 0.31 1.00 1.00 1.00 0.85
Claude3 H 0.68 0.29 0.57 0.65 0.40 0.53 0.33
GPT-4o 0.96 0.31 0.77 0.83 0.73 0.70 0.63
@@
LLaVA v1.6 0.39 0.27 0.32 0.45 0.45 0.42 0.45
Phi-4 M.I. 0.62 0.29 0.45 0.70 0.82 0.68 0.70
Idefics 2 0.49 0.25 0.29 0.35 0.40 0.40 0.50
SmolVLM 0.26 0.34 0.28 0.40 0.48 0.44 0.60
QwenVL 0.68 0.27 0.61 0.70 0.50 0.60 0.40
Llama 3.2 0.66 0.24 0.27 0.45 0.63 0.43 0.30
Human 0.81 0.95 0.83 0.93 0.85 0.95 0.90
I.1 Temporal Matching Task
Hourly Time Series
Start time: 2022-07-12 22:00:00
[19.59, 19.48, 20.19, 18.95, 19.53, 19.41, 19.53, 18.54, 19.09, 16.71, 18.26, 19.19, 19.01, 19.38, 19.60, 19.63, 17.98, 17.71, 18.03, 18.43, 18.51, 18.30, 18.40, 18.16, 18.16, 18.42, 19.49, 17.91, 19.19, 18.20, 18.80, 18.20, 18.26, 17.87, 18.03, 18.29, 17.84, 18.73, 18.88, 18.27, 17.97, 18.09, 17.96, 17.37, 18.05, 18.42, 18.00, 18.73, 17.54, 18.56, 17.58, 17.73, 18.42, 18.02, 17.94, 17.75, 17.72, 19.24, 18.87, 18.74, 18.01, 17.88]
Early Bucket  Query: 2022-07-13 15:00:00 Options: A. 17.71 B. 19.38 C. 17.98 D. 18.16 Answer: A Mid-1 Bucket Query: 2022-07-14 00:00:$0$ Options: A. 18.80 B. 18.26 C. 18.40 D. 19.49 Answer: D Mid-2 Bucket Query: 2022-07-14 04:00:00 Options: A. 17.84 B. 18.80 C. 18.03 D. 18.42 Answer: B Mid-3 Bucket Query: 2022-07-14 18:00:$0$ Options: A. 18.56 B. 18.42 C. 18.05 D. 18.73 Answer: C Late Bucket Query: 2022-07-15 04:00:$0$ Options: A. 17.94 B. 18.01 C. 18.56 D. 18.02 Answer: A
Figure 11: Illustration of temporal buckets used in our diagnostic MCQ task. All questions query the same hourly time series, with increasing temporal distance from the start timestamp, isolating temporal indexing difficulty.
Distractor Generation for Temporal Matching
Distractor options are generated using a temporal locality–driven sampling strategy. Given a time series $\mathbf{x}=\{x_{1},\dots,x_{N}\}$ and a correct index $i^{\ast}$, we first collect candidate values from a symmetric temporal window centered at $i^{\ast}$. If the number of unique candidates is insufficient, the window is expanded to a larger radius, with a final fallback to sampling from the entire series. This ensures distractors are temporally plausible while avoiding trivial elimination strategies.
Algorithm 1 Temporal Locality–Driven Distractor Sampling (TLDS)
0: Time series $\mathbf{x}$, correct index $i^{\ast}$, window size $w$, maximum window $W$, number of options $K$
0:K1K-1 distractor values
1: Initialize empty candidate set $\mathcal{C}$
2:for o=wo=-w to $w$ do
3:  $i\leftarrow i^{\ast}+o$
4:  if $0\leq i<N$ and $i\neq i^{\ast}$ then
5:   if $x_{i}\neq x_{i^{\ast}}$ then
6:    Add $x_{i}$ to $\mathcal{C}$
7:   end if
8:  end if
9:end for
10:if $|\mathcal{C}|<K-1$ then
11:  Repeat the above procedure using window size $W$
12:end if
13:if $|\mathcal{C}|<K-1$ then
14:  $\mathcal{C}\leftarrow\{x_{1},\dots,x_{N}\}\setminus\{x_{i^{\ast}}\}$
15:end if
16: Sample K1K-1 distinct values uniformly from $\mathcal{C}$
17:return sampled distractors
Index Lookup Baseline.
We implement a strict index-based lookup baseline that deterministically retrieves the ground-truth value for a queried timestamp when exact temporal alignment is possible. Given a start time and query time, the baseline computes the hour offset using strict datetime parsing and requires the offset to be an exact integer. If the resulting index is within bounds, the corresponding time-series value is retrieved and matched against the multiple-choice options using exact numeric equality, with no rounding or tolerance. The baseline returns an answer only when a unique exact match exists; otherwise, it abstains. This baseline serves as a correctness upper bound for the temporal matching task and isolates model failures from ambiguities in data representation or parsing.
Algorithm 2 Temporal Index Lookup Baseline
0: Time series $\mathbf{x}=[x_{0},\dots,x_{N-1}]$, start time $t_{0}$, query time $t_{q}$, option set $\mathcal{O}$
0: Predicted option label or failure
1: Parse $t_{0}$ and $t_{q}$ using ISO datetime parsing
2: Compute time offset $\Delta t\leftarrow t_{q}-t_{0}$
3: Convert offset to hours: $h\leftarrow\Delta t/3600$
4:if $h$ is not an integer then
5:  return failure
6:end if
[…]
8:if $i<0$ OR $i\geq|\mathbf{x}|$ then
9:  return failure
10:end if
[…]
12: Parse option values using numeric parsing
[…]
14:if $|\mathcal{M}|=1$ then
15:  return the unique label in $\mathcal{M}$
16:else
17:  return failure
18:end if
Table 20: MCQ accuracy on short mid2 questions when the time series is provided as numeric text, as a plotted image, or as both image and text. Vision input improves performance for most models, but does not resolve failures in exact temporal indexing.
Model Text Only Image Only Image+Text
Gemini-Flash 46.7 57.3 34.7
Claude-3 Haiku 33.3 42.7 38.7
GPT-4o 46.7 60.0 57.3
GPT-5.1 (openai2025gpt51) 46.7 58.7 54.7
InternVL3-14B 50.7 60.0 74.7
InternVL3-38B 33.3 46.7 76.0
LLaMA-90B 44.0 44.0 38.7
Gemma 57.3 58.7 58.7
Modality Ablations.
Table 20 compares model accuracy on mid2 Temporal Matching questions when the same underlying time series is provided as numeric text, as a plotted image, or as both image and text. Image-only input yields higher accuracy than text-only input for most models and often reaches the highest performance for a given model, suggesting that visual structure strongly supports coarse temporal localization. However, adding numeric text alongside the image frequently degrades performance, indicating that many models struggle to jointly integrate visual and textual representations of the same time series. This multimodal interference is evident for models such as Gemini 2.0 Flash, GPT-4o, and LLaMA-90B, where image-only accuracy exceeds image-and-text performance. Notably, only the newer InternVL3 (zhu2025internvl3exploringadvancedtraining) models, particularly the larger InternVL variants, consistently benefit from multimodal input and achieve the highest accuracies in the combined setting. These results suggest that while vision alone can aid approximate temporal reasoning, effective fusion of redundant visual and numeric signals remains a challenge for most models, with successful multimodal integration emerging only in more recent architectures.
I.2 Sample Q&A Questions
@@
We provide examples of Q&A questions in Figures 121316151920, and 14, covering one example per question type.
@@
Here is a time series caption:
Answer
"answer": "D"
Figure 14: Example of a time series matching question.
Question
@@
You are given an hourly time series.
Time series: [19.59, 19.48, 20.19, 18.95, 19.53, 19.41, 19.53, 18.54, 19.09, 16.71, 18.26, 19.19, 19.01, 19.38, 19.6, 19.63, 17.98, 17.71, 18.03, 18.43, 18.51, 18.3, 18.4, 18.16, 18.16, 18.42, 19.49, 17.91, 19.19, 18.2, 18.8, 18.2, 18.26, 17.87, 18.03, 18.29, 17.84, 18.73, 18.88, 18.27, 17.97, 18.09, 17.96, 17.37, 18.05, 18.42, 18.0, 18.73, 17.54, 18.56, 17.58, 17.73, 18.42, 18.02, 17.94, 17.75, 17.72, 19.24, 18.87, 18.74, 18.01, 17.88]
The series starts at: 2022-07-12 22:00:00
Question: What is the value of the time series at the following time? 2022-07-13 15:00:00
Options:
A. 17.71  B. 19.38  C. 17.98  D. 18.16
You MUST answer using exactly one capital letter. Final answer (one letter only):
Answer
A
Figure 17: Example of a time series temporal matching question.
QuestionRefer to caption You are given an hourly time series shown as a plot. Only the starting time is labeled on the x-axis. Each point represents one hour. Start time: 2022-07-14 01:00:00 Question: What is the value of the time series at the following time? 2022-07-14 04:00:00 Options: A. 18.26  B. 18.8  C. 19.19  D. 18.2 Answer with exactly one capital letter from {A, B, C, D}. Do not explain anything. Just answer in one capital letter. Answer: Answer
B
Figure 18: Example of a plot-based temporal matching question.
Question
@@
I.3 Qwen-Based Filtering
@@
I.4 Distractor Generation in Q&A Tasks
@@
TS Matching Cross-domain 0.9803
TS Matching Same-domain 0.9586
TS Matching Artificially Perturbed 0.6864
@@
Appendix J Inference Examples
J.1 Numeric Time Series Ablation
@@
J.2 Statistical Inference Failure & Success
@@
Appendix K Editing and Review
K.1 Edits made in the captions
@@
K.2 Semantic Drift Between Unedited and Edited Captions
In this section, we show how much the captions were edited quantitatively. Table 23 quantifies the extent of human edits across domains using word-level divergence (WER), token-level change, and numeric overlap. The magnitude of rewriting varies substantially by domain. For example, border crossing, diet, and Walmart captions exhibit very large edits, with mean WER values of 0.944, 0.885, and 0.939 respectively, alongside sizeable average token increases of +28.04, +26.88, and +21.24, indicating extensive rewriting and content addition.
Other domains reflect more moderate editing. Agriculture and CO2 show mean WER values of 0.663 and 0.733, with average token increases of +7.63 and +18.76, suggesting partial restructuring rather than complete rewrites. Further, domains such as COVID, road injuries, and online retail exhibit lower rewrite magnitudes, with WER values of 0.232, 0.132, and 0.396 respectively, and small or negative mean token deltas, indicating that many edits involve localized refinements or concision rather than expansion.
Across domains, numeric Jaccard scores remain relatively high in several cases (e.g., 0.956 for demography and 0.892 for agriculture), showing that even when captions are substantially rewritten, numeric content is often preserved. Overall, human editing spans a wide spectrum of intervention levels, ranging from targeted phrasing adjustments to near-complete caption rewrites, depending on domain characteristics and the needs of individual captions.
Table 23: Domain-wise mean rewrite magnitude (WER), lexical overlap (ROUGE-L), numeric faithfulness (Numeric Jaccard), and caption length change ($\Delta$ Tokens) between unedited and human-edited captions.
Domain WER ROUGE-L Numeric Jaccard $\Delta$ Tokens
Agriculture 0.663 0.541 0.892 7.63
Border Crossing 0.944 0.287 0.509 28.04
CO2 0.733 0.452 0.674 18.76
COVID 0.232 0.859 0.846 -2.95
Crime 0.593 0.555 0.854 0.88
Demography 0.504 0.637 0.956 4.67
Diet 0.885 0.365 0.618 26.88
Online Retail 0.396 0.730 0.816 -10.76
Road Injuries 0.132 0.921 0.862 2.68
Walmart 0.939 0.288 0.329 21.24
Average 0.602 0.564 0.736 9.71
Figure 24: Lexical overlap versus rewrite magnitude across domains. Each point corresponds to a single caption pair, plotted by word error rate (WER) between the unedited and human-edited captions (x-axis) and ROUGE-L overlap (y-axis). Colors denote different domains. We see a strong inverse relationship between rewrite magnitude and lexical overlap: captions with low WER retain high ROUGE-L scores, while captions undergoing substantial rewriting (WER approaching 1.0) exhibit markedly lower lexical overlap. This consistent trend across domains indicates that increased human intervention is primarily expressed through lexical and structural rephrasing. This plot also highlights how human edits progressively reduce surface-level overlap as rewrite magnitude increases, reflecting varying degrees of paraphrasing and content restructuring.
Figure 25: Numeric preservation versus rewrite magnitude across domains. Each point represents a single caption pair, plotted by word error rate (WER) between the unedited and human-edited captions (x-axis) and numeric Jaccard similarity (y-axis). Colors indicate different domains. The figure shows that captions span a wide range of rewrite magnitudes, from minor edits (low WER) to near-complete rewrites (WER close to 1.0). Despite substantial rewriting in many cases, numeric overlap often remains high, with numerous samples exhibiting numeric Jaccard scores above 0.8 even at high WER values. At the same time, certain domains display greater dispersion at high rewrite magnitudes, reflecting more extensive numeric restructuring when captions are heavily rewritten.
K.3 Interface
@@
Appendix L Template-based Prompts
@@
L.1 Ground-Truth Synthetic Caption Generation Prompt
@@
L.2 Baseline Caption Generation Prompt
@@
Here is a time series about <sampling frequency> <measure> in the Indian city of <city name>.
[…]
Describe this time series following these rules:
1. specify and include all contextual data that is needed.
2. make statistical comparisons (max, min, mean, standard deviation) by providing numerical values.
3. comment on patterns, shapes, and trends of the time series.
4. explain how the period compares to historical or global norms.
5. describe the nature of the movements, patterns, and anomalies if any.
6. walk through the timeline and describe sub-periods, phases, and periodic patterns when present.
7. Report dates when events happened.
**Not all information for the above rules might be present or needed**, focus on information that is actually evident in the data inputs.
Here is some historical context:
- all-time maximum: <all-time maximum>
- all-time minimum: <all-time minimum>
- all-time average value until today: <all-time average value until today>
- all-time standard deviation until today: <all-time standard deviation until today>
Avoid repetitive phrasing and overused expressions.
Answer in a single paragraph of four to five sentences at most, without bullet points or any formatting.
L.3 Caption Paraphrasing Prompt
@@
The full TSC prompt from L.2 is injected as the caption_prompt string.
L.5 Semantic Perturbation Prompt
@@
L.6 Numeric Perturbation Prompt